OpenStudy (anonymous):

All of the following are equivalent to 3:8=x:32 except 8:3 = 32:x 3:x = 8:32 x:8 = 3:32 8:32 = 3:x

OpenStudy (callisto):

3:8=x:32 => \[\frac{3}{8} = \frac{x}{32}\] (a) 8:3 = 32:x \[\frac{8}{3} = \frac{32}{x}\] Take reciprocal for both sides \[\frac{3}{8} = \frac{x}{32}\] (b) 3:x = 8:32 \[\frac{3}{x} = \frac{8}{32}\]\[\frac{3}{8} = \frac{x}{32}\] (c) x:8 = 3:32 \[\frac{x}{8} = \frac{3}{32}\]\[32x =8(3)\] (d) 8:32 = 3:x \[\frac{8}{32} = \frac{3}{x}\] \[\frac{x}{32} = \frac{3}{8}\] What's the answer